# Too Good to be Bad: On the Failure of LLMs to Role-Play Villains

### v1

@@ 1 Introduction
Our findings provide the first large-scale empirical evidence that LLMs systematically struggle with antagonistic role-play. We observe a consistent, monotonic decline in performance as a character’s morality decreases, with average fidelity scores dropping from 3.21 for moral paragons to 2.61 for villains. The most significant performance degradation occurs at the boundary between flawed-but-good (Level 2) and egoistic (Level 3) characters, suggesting that the inability to simulate self-serving behavior is a primary obstacle. A fine-grained analysis reveals that models are most heavily penalized for failing to portray negative traits like “Manipulative”, “Deceitful”, and “Cruel”, which directly conflict with the principles of helpful and honest AI. Furthermore, we find that a model’s general conversational ability, as measured by leaderboards like the Arena, is a poor predictor of its villain role-playing skill. This is particularly evident for highly-aligned models, which show a disproportionate drop in performance when tasked with portraying villainy.
@@ 3.1 Main Results
Averaging across all evaluated models, performance drops from 3.21 (Level 1) and 3.13 (Level 2) to 2.71 (Level 3) and 2.61 (Level 4). The largest decline occurs between Level 2 and Level 3 (-0.42), while the drop from Level 3 to Level 4 (-0.10) is comparatively smaller. This pattern directly supports our central claim that antagonistic role-play is systematically harder: models are relatively strong on benevolent or mildly flawed personas but falter when asked to embody self-serving and overtly villainous characters. These aggregate numbers confirm the contribution that role-playing quality degrades as morality decreases, with the pivotal difficulty appearing at the egoist boundary.
@@
Examining per-model deltas from Level 2 to Level 3 shows consistent, substantial drops: qwen3-max (-0.65), grok-4 (-0.56), claude-sonnet-4.5 (-0.48), deepseek-v3 (-0.46), claude-opus-4.1 (-0.45), o3 (-0.42), and chatgpt-4o-latest (-0.41). Even the better-performing families (gemini-2.5-pro: -0.30; deepseek-v3.1: -0.28; deepseek-v3.1-thinking: -0.24) show clear degradation. This indicates a universal modeling challenge at the onset of egoistic, manipulative behavior, aligning with our hypothesis that alignment and training biases toward prosocial helpfulness suppress authentic simulation of self-serving personas.
@@ 4.1 Villain RolePlay Leaderboard
chatgpt-4o-latest 7 2.65 2 1440
deepseek-R1 8 2.62 11 1417
claude-sonnet-4.5 9 2.56 2 1438
glm-4.5 10 2.55 18 1406
claude-sonnet-4.5-thinking 11 2.54 1 1445
grok-4 12 2.54 12 1413
claude-opus-4.1-thinking 13 2.53 1 1447
grok-4-fast 14 2.50 11 1420
claude-opus-4.1 15 2.48 2 1437
deepseek-v3 16 2.41 36 1391
qwen3-max 17 2.33 10 1423
@@
Our findings, summarized in the Villain RolePlay (VRP) leaderboard in Table 7, reveal a significant misalignment between a model’s general aptitude (Arena Rank) and its specialized ability to portray villains. For example, glm-4.6, which ranks first in villain role-play, is only tenth in the general Arena. Conversely, top-tier Arena models like gemini-2.5-pro (Arena Rank 1, VRP Rank 4) and claude-opus-4.1-thinking (Arena Rank 1, VRP Rank 13) demonstrate a notable drop in relative performance. This discrepancy strongly supports our central claim that the skills required for helpful, harmless conversation are distinct from, and may even conflict with, those needed for authentic antagonistic role-play.
@@
The trend is most pronounced for models renowned for their strong safety alignment, such as the Claude family. Despite their top rankings in general-purpose benchmarks, claude-sonnet-4.5 and claude-opus-4.1 place 9th and 15th respectively on our VRP leaderboard. This steep decline provides compelling evidence for our hypothesis that robust safety guardrails, while crucial, can systematically hinder a model’s ability to simulate malevolent or manipulative personas, even within a clearly demarcated fictional context. The superior performance of models like glm-4.6 may indicate that their alignment strategies are more context-aware, allowing for greater fidelity in character simulation.

### v2

@@ 1 Introduction
Our findings provide the first large-scale empirical evidence that LLMs systematically struggle with antagonistic role-play. We observe a consistent, monotonic decline in performance as a character’s morality decreases, with average fidelity scores dropping from 3.21 for moral paragons to 2.62 for villains. The most significant performance degradation occurs at the boundary between flawed-but-good (Level 2) and egoistic (Level 3) characters, suggesting that the inability to simulate self-serving behavior is a primary obstacle. A fine-grained analysis reveals that models are most heavily penalized for failing to portray negative traits like “Manipulative”, “Deceitful”, and “Cruel”, which directly conflict with the principles of helpful and honest AI. Furthermore, we find that a model’s general conversational ability, as measured by leaderboards like the Arena, is a poor predictor of its villain role-playing skill. This is particularly evident for highly-aligned models, which show a disproportionate drop in performance when tasked with portraying villainy.
@@ 3.1 Main Results
Averaging across all evaluated models, performance drops from 3.21 (Level 1) and 3.14 (Level 2) to 2.71 (Level 3) and 2.62 (Level 4). The largest decline occurs between Level 2 and Level 3 (-0.43), while the drop from Level 3 to Level 4 (-0.09) is comparatively smaller. This pattern directly supports our central claim that antagonistic role-play is systematically harder: models are relatively strong on benevolent or mildly flawed personas but falter when asked to embody self-serving and overtly villainous characters. These aggregate numbers confirm the contribution that role-playing quality degrades as morality decreases, with the pivotal difficulty appearing at the egoist boundary.
@@
Examining per-model deltas from Level 2 to Level 3 shows consistent, substantial drops: qwen3-max (-0.65), grok-4 (-0.56), claude-sonnet-4.5 (-0.48), hunyuan-turbos (-0.48), deepseek-v3 (-0.46), claude-opus-4.1 (-0.45), o3 (-0.42), and chatgpt-4o-latest (-0.41). Even the better-performing families (gemini-2.5-pro: -0.30; deepseek-v3.1: -0.28; deepseek-v3.1-thinking: -0.24) show clear degradation. This indicates a universal modeling challenge at the onset of egoistic, manipulative behavior, aligning with our hypothesis that alignment and training biases toward prosocial helpfulness suppress authentic simulation of self-serving personas.
@@ 4.1 Villain RolePlay Leaderboard
hunyuan-turbos 7 2.66 49 1380
chatgpt-4o-latest 8 2.65 2 1440
deepseek-R1 9 2.62 11 1417
claude-sonnet-4.5 10 2.56 2 1438
glm-4.5 11 2.55 18 1406
claude-sonnet-4.5-thinking 12 2.54 1 1445
grok-4 13 2.54 12 1413
claude-opus-4.1-thinking 14 2.53 1 1447
grok-4-fast 15 2.50 11 1420
claude-opus-4.1 16 2.48 2 1437
deepseek-v3 17 2.41 36 1391
qwen3-max 18 2.33 10 1423
@@
Our findings, summarized in the Villain RolePlay (VRP) leaderboard in Table 7, reveal a significant misalignment between a model’s general aptitude (Arena Rank) and its specialized ability to portray villains. For example, glm-4.6, which ranks first in villain role-play, is only tenth in the general Arena. Conversely, top-tier Arena models like gemini-2.5-pro (Arena Rank 1, VRP Rank 4) and claude-opus-4.1-thinking (Arena Rank 1, VRP Rank 14) demonstrate a notable drop in relative performance. This discrepancy strongly supports our central claim that the skills required for helpful, harmless conversation are distinct from, and may even conflict with, those needed for authentic antagonistic role-play.
@@
The trend is most pronounced for models renowned for their strong safety alignment, such as the Claude family. Despite their top rankings in general-purpose benchmarks, claude-sonnet-4.5 and claude-opus-4.1 place 10th and 16th respectively on our VRP leaderboard. This steep decline provides compelling evidence for our hypothesis that robust safety guardrails, while crucial, can systematically hinder a model’s ability to simulate malevolent or manipulative personas, even within a clearly demarcated fictional context. The superior performance of models like glm-4.6 may indicate that their alignment strategies are more context-aware, allowing for greater fidelity in character simulation.